OpenStudy (anonymous):

r2-2rs+s2=4, (r-s)^6 ?

OpenStudy (dumbcow):

did you mean (r-s)^2 ?

OpenStudy (anonymous):

no its (r-s)^6

OpenStudy (dumbcow):

?? i don't understand the problem, are we solving for "r"

OpenStudy (anonymous):

we are solving for (r-s)^6 that what will it be equalled to

OpenStudy (anonymous):

Here is another.. if the ratio of area of 2 squares is 2:1 then what will be the ratio of circumference of circles?

OpenStudy (dumbcow):

oh ok by factoring left side (r-s)^2 = 4 so (r-s)^6 = 4^3 = 64

OpenStudy (anonymous):

thanks. Solve the other other plz

OpenStudy (dumbcow):

what .... how are the squares related to the circles?

OpenStudy (anonymous):

i think it we have to find radius for both squares and then put in equations for the circumference